OpenStudy (anonymous):

what is the derivative of 8sin(4pix)

OpenStudy (zarkon):

you know the derivative of \(\sin(x)\)?

OpenStudy (anonymous):

-cos(x)

OpenStudy (zarkon):

you know the chain rule?

OpenStudy (anonymous):

is it (f(g(x))) = f'(g(x))g'(x)

OpenStudy (zarkon):

put your two answers together :)

OpenStudy (anonymous):

-8cos(4pix)

OpenStudy (anonymous):

(4pix)

OpenStudy (fretje):

is 8.cos(4.pi.x) if derived to 4.pi.x if derived to x: 8.4.pi.cos (4.pi.x)

OpenStudy (zarkon):

close...you need to multiply by the derivative of the inside

OpenStudy (zarkon):

yes...plus \[\frac{d}{dx}\sin(x)=\cos(x)\]

OpenStudy (anonymous):

-32cos(x)

OpenStudy (zarkon):

and don't forget the \(\pi\)

OpenStudy (anonymous):

so the answer is -32cos(pix)

OpenStudy (zarkon):

\[32\pi\cos(4\pi x)\]

OpenStudy (anonymous):

8sin(4πx) ( sin u )' = u' cos u u = 4πx --> u' = 4π -> 8 * 4π cos (4πx) = 32 π cos (4πx)

OpenStudy (anonymous):

Remember the sign is positive :)